OpenStudy (anonymous):

WILL MEDAL Simon invested $1,550 at 6.5% simple interest. He earned $302.25 in interest after t years. What is the value of t? ()1 ()9 ()6 ()3

OpenStudy (anonymous):

PLZ HELP

OpenStudy (kropot72):

The interest after once year is 1,555 * 0.065. Now we can set up an equation for the interest after t years: \[\large 302.25=t(1550\times0.065)\ .....(1)\] Now you just need to solve equation (1) to find the value of t.

OpenStudy (anonymous):

how

OpenStudy (kropot72):

You just divide both sides of equation (1) by (1550 * 0.065): \[\large t=\frac{302.25}{(1550\times0.065)}=you\ can\ calculate\]